The sample BOD5 measured was 15mg/L and the rate constant was 0.1/day, what would be
the ultimate BODult (oxygen required to stabilize the total quantity of biologically oxidizable
organic matter present) of this sample. (Hint: L = Lo 10-kt)

thanks
Answer #1

sols ginen Bons = 15 male = L K = Ool /day t = 5 days Lo= ultimate BOD 4 = Lo lokt { given in hint} BOD=> L = lo-Lť. L=Lo-to